# Descriptive Caption Enhancement with Visual Specialists for Multimodal Perception

### v2

@@ 3.2 Object Relation
LLaVA-NeXT [24] Vicuna-7B 81.8 74.4 44.3 64.2 64.9 70.1 66.6
@@
LLaVA-NeXT(Ours) Vicuna-7B 82.4 78.8 57.2 65.2 64.8 71.2 71.2
3.3 Captioning with Attributes
Region captioning. We use a large language model (Qwen-72B) to integrate the object attributes with the caption from a large multimodal model (InternVL2-26B). For example, in cases where the fine-grained model for “animals” identifies a specific class label like “$\{animal\_name\}$, we employ the prompt $\{cat\_name\}$ exists in the region and $\{animal\_name\}$ is the $\{cat\_name\}$’s subclass; use $\{animal\_name\}$ in the caption; otherwise, do not mention $\{animal\_name\}$ to guide the LLM in integrating this specific label into the caption. In this process, $\{cat\_name\}$ represents the coarse-grained label provided by the detection model. The detailed prompt engineering process is illustrated in Figure 4.
Image captioning. We combine object relation attributes, and region location information (for object grounding) and captions to get improved image captions. We use LLM for the combination. The prompt is given in Figure 5. For example, to describe the 3D relative positional relationship between $\{bbox\_0\}$ and $\{bbox\_1\}$, we use the prompt: ”Relative to the camera, the $\{category\_0\}$ in $\{bbox\_0\}$ in the image is $\{3d\_relation\}$$\{category\_1\}$ in $\{bbox\_1\}$, where $\{3d\_relation\}$ can be ”in front of” or ”behind of”. We then use a large language model (Qwen2-72B-AWQ) to integrate this information with the detailed region caption and the image reference caption.
3.4 Analysis
Dataset Description. We leverages DCE to generate more detailed annotations for publicly available image datasets, resulting in two enhanced datasets: (1) DCE-1M, which provides dense annotations for 1 million diverse images from Densefusion [26], covering a wide range of objects and scenes, and (2) DCE-118K, which includes refined annotations for 118,000 complex scene images from the COCO dataset [30]. We provide a detailed analysis of the DCE-annotated image captions in the Appendix.
Attribute richness. We randomly select 100 images from the DCE-118K and invite five independent evaluators to assess the captions for each image. The evaluators analyzed and recorded the occurrence of the attributes within the captions, such as spatial relations, HOI, and OCR (optical character recognition) details.
Comparison. We conducted a comparison of human annotations, internVL2-26B, LLaVA-NeXT-34B, and DCE-annotated 118K datasets on both LLaVA-v1.5 and LLaVA-NeXT. The results as shown in Figure 2, the captions annotated by DCE significantly improve the performance of LLaVA-v1.5 and LLaVA-NeXT. Compared to human and generalist multimodal model annotations, DCE-annotated captions are richer and more detailed, offering deeper context and capturing finer nuances. This enhanced descriptive quality leads to better performance in downstream tasks.
4 Experiments
In this section, we first present the implementation details, then compare the DCE-1M-trained model with state-of-the-art LMMs across multiple benchmarks. Finally, we perform ablation studies on the DCE-118K-trained model to validate the effectiveness of DCE.
@@ 4 Experiments
Method Vision Encoder Language Model MMBench-CN MMBench MM-Vet SEEDI SEED-Bench MMMU POPE
@@
LLaVA-NeXT(Ours) ViT-L (0.3B) Vicuna-7B 61.7 69.3 40.1 72.2 65.7 36.0 87.0
@@
Model and Training Set. We conduct experiments on LLaVA-v1.5 [31] and LLaVA-NeXT [32] to demonstrate the effectiveness of DCE. Specifically, we using CLIP-L [42] as the visual encoder and Vicuna-v1.5 [12] as the large language model. We adopt a two-stage training strategy: (1) Pre-Training Stage. We train only the projector for initial alignment. Then, following SharGPT4V [8], we set the last 12 layers of the visual encoder in LLaVA-v1.5 as trainable and make the entire LLaVA-NeXT model trainable, following [23], to further enhance perceptual capabilities. (2) Instruction Tuning Stage. We use the open-source LLaVA-mix-665K and LLaVA-NeXT-data to respectively train the LLaVA-v1.5 and LLaVA-NeXT models. The detailed training procedure is provided in the supplementary material.
@@
Comparison Method. We compare LLaVA-v1.5 and LLaVA-NeXT models trained on DCE-1M against current SOTA MLLMs. Additionally, we evaluate the performance of various MLLMs in generating image captions, using the advanced models InternVL2-26B [11] and LLaVA-NeXT-34B [24] to generate competitive captions on the DCE-118K images.
@@ 4.2 Main Results
VQA Benchmarks. The results on six common visual question answering (VQA) datasets are presented in Table 3. It is clear that LLaVA-v1.5 and LLaVA-NeXT, trained with DCE-1M, achieve state-of-the-art performance in both low-resolution and high-resolution settings. Compared to the baseline LLaVA-v1.5 [31], our model excels across all VQA benchmarks, demonstrating that high-quality image captions significantly enhance model performance. This highlights the crucial role of detailed and accurate captions in improving visual understanding for VQA tasks. Furthermore, when compared to the baseline LLaVA-NeXT [24], this improvement remains consistent, suggesting that the impact of high-quality captions is not dependent on the model variation. Additionally, compared to models like ShareGPT-4V, our model demonstrates superior performance across most VQA benchmarks. This improvement indicates that the captions generated by our DCE method provide richer and more comprehensive information.
@@
      Annotation       OKVQA       GQA       ScienceQA       TextVQA       MMBench       MM-Vet       SEED-Bench
@@
      + DCE       56.7       65.2       72.0       62.0       68.5       37.8       65.0
Figure 6: Visualization of DCE’s Attribute Fusion: DCE combines object and relational attributes to generate detailed and comprehensive captions.
The model trained on DCE-1M demonstrates exceptional performance on datasets such as VQAv2 [15] and Ai2D [19], highlighting that integrating detection models into DCE significantly enriches the diversity of objects in the generated captions, thereby boosting the model’s performance on object recognition benchmarks. Furthermore, incorporating relational attributes into DCE enriches the captions with detailed inter-object relationships, enhancing the model’s ability to capture and understand complex relationships, which further improves its performance on visual reasoning benchmarks such as GQA [17]. However, performance on tasks like TextVQA [47] is hindered by limitations in the open-source OCR model and the threshold settings; a high threshold restricts the model’s ability to capture finer textual details.
Large Multi-Modal Benchmarks. We further conduct the evaluation on five challenging large multi-modal benchmarks. The experimental results are shown in Table 4. It can be seen that both LLaVA-v1.5 and LLaVA-NeXT trained with DCE-1M achieve competitive performance on more complex LMM benchmarks, demonstrating that the improvements brought by DCE-1M are comprehensive. Our model outperforms both LLaVA-v1.5 [31] and LLaVA-NeXT [32] across all LMM benchmarks, demonstrating that high-quality image captions during pretraining significantly enhance model performance, even without altering the supervised fine-tuning (SFT) data. Compared to other image captioning methods, such as ShareGPT-4V [8], DCE-generated captions provide richer and more comprehensive scene information, significantly boosting model performance across most LMM benchmarks. However, due to the lack of Chinese data in DCE-1M, the model performs poorly on MMBench-CN [34]. This highlights the need for multilingual image captioning, which will be an area for future improvement in DCE. Additionally, the detection model in DCE introduces some noise, which may interfere with the model’s ability to accurately capture objects, leading to decreased performance on tasks like POPE [27]. Therefore, reducing this noise will be a key focus for future improvements in DCE.
@@ 4.3 Ablation Study
Comparing different annotation methods. We compared different image annotation methods, including human annotations, GenerateList LMM annotations, and our DCE. Specifically, we annotated 118K COCO images and conducted comparisons on LLaVA-v1.5 and LLaVA-NeXT. The experimental results are shown in Table 5. We found that the image captions generated by DCE improve LMM performance on downstream tasks more effectively than other annotation methods. Notably, compared to captions annotated by internVL2, DCE’s inclusion of object attributes significantly improves model performance on OKVQA and TextVQA tasks. Relational attributes enhance the model’s understanding of multi-object relationships, leading to a notable increase in GQA performance. Furthermore, superior results on complex LMM benchmarks like MM-Vet, MMBench, and SEED highlight that DCE-generated captions provide rich and comprehensive scene information.
Case Study. Figure 6 presents example captions generated by the DCE engine and the general MLLM LLaVA-NeXT 34B. It is evident that visual specialists within DCE capture detailed object and relational attributes, resulting in richer and more descriptive captions. For instance, in Figure 6(a), the OCR information highlights the precision of OCR specialist model, while the fine-grained model’s identification of specific aircraft types further enhances caption informativeness. Additionally, the relational attributes significantly enrich description by providing detailed spatial relationships between objects, underscoring the advantages of DCE in capturing comprehensive scene details. More visualization results are provided in the Appendix.
@@ 5 Conclusion
In this paper, we propose a new image captioning engine, DCE , which utilizes off-the-shelf visual specialists to enhance the quality and detail of image captions. Unlike existing methods that primarily rely on LMM models or human annotation, DCE leverages visual specialists to simulate human perceptual abilities, while using LLMs to emulate cognitive processes. This dual approach enables DCE to generate captions that are both visually detailed and contextually aware. Through extensive experiments, we demonstrate that incorporating these visual specialists leads to improved model performance across various visual understanding and reasoning tasks, particularly those that depend on accurate attribute and relationship recognition. Our approach highlights the potential of leveraging specialized visual features for enhancing multimodal representations and provides a flexible framework for future integration of additional visual expertise. We will release the DCE source code and pipeline to facilitate further research, enabling the community to easily integrate other visual specialists and extend the capabilities of multimodal models.

### v3

@@ 4 Experiments
LLaVA-NeXT [27] Vicuna-7B 81.8 74.4 54.3 64.2 64.9 70.1 66.6
LLaVA-NeXT+ReCap [26] Vicuna-7B - 75.3 44.3 65.0 - 71.0 66.9
@@
lightgray black LLaVA-NeXT(Ours) Vicuna-7B 82.4 78.8 57.2 65.2 64.8 71.2 71.2
In this section, we first present the implementation details, then compare the EDC-1M-trained model with state-of-the-art LMMs across multiple benchmarks. Finally, we perform ablation studies on the EDC-118K-trained model to validate the effectiveness of EDC.
@@
Method Vision Encoder Language Model MMBench-CN MMBench $M$-Vet SEEDI SEED-Bench MMMU POPE
@@
lightgray black LLaVA-NeXT(Ours) ViT-L (0.3B) Vicuna-7B 61.7 69.3 40.1 72.2 65.7 36.0 87.0
@@
Model and Training Set. We conduct experiments on LLaVA-v1.5 [34] and LLaVA-NeXT [35] to demonstrate the effectiveness of EDC. Specifically, we using CLIP-L [48] as the visual encoder and Vicuna-v1.5 [11] as the large language model. We adopt a two-stage training strategy: (1) Pre-Training Stage. We train only the projector for initial alignment. Then, following SharGPT4V [7], we set the last 12 layers of the visual encoder in LLaVA-v1.5 as trainable and make the entire LLaVA-NeXT model trainable, following [26], to further enhance perceptual capabilities. (2) Instruction Tuning Stage. We use the open-source LLaVA-mix-665K and LLaVA-NeXT-data to respectively train the LLaVA-v1.5 and LLaVA-NeXT models. The detailed training procedure is provided in the supplementary material.
@@
Comparison Method. We compare LLaVA-v1.5 and LLaVA-NeXT models trained on EDC-1M against current SOTA MLLMs. Additionally, we evaluate the performance of various MLLMs in generating image captions, using the advanced models InternVL2-26B [10] and LLaVA-NeXT-34B [27] to generate competitive captions on the EDC-118K images.
4.2 Comparison with the State-of-the-art
To validate the effectiveness of EDC captions, we compare them with several state-of-the-art image captioning methods on the LLaVA and LLaVA-Next benchmarks. These methods include ShareGPT4V [7], DenseFusion [29], ReCap [26], DCI [62], DOCCI [43], and IT [46]. As shown in Table 3, models trained with EDC captions achieve better results on most evaluation metrics.
This demonstrates that EDC captions enhance LMM training by providing more comprehensive image descriptions, capturing finer-grained visual attributes and richer contextual information compared to existing methods. This performance gain can be attributed to the design of EDC, which integrates fine-grained object details, 3D spatial relations, and grounded region-level information via structured prompt engineering, offering more informative and aligned supervision signals.
@@ 4.3 Main Results
VQA Benchmarks. The results on six common visual question answering (VQA) datasets are presented in Table 4. It is clear that LLaVA-v1.5 and LLaVA-NeXT, trained with EDC-1M, achieve state-of-the-art performance in both low-resolution and high-resolution settings. Compared to the baseline LLaVA-v1.5 [34], our model excels across all VQA benchmarks, demonstrating that high-quality image captions significantly enhance model performance. This highlights the crucial role of detailed and accurate captions in improving visual understanding for VQA tasks. Furthermore, when compared to the baseline LLaVA-NeXT [27], this improvement remains consistent, suggesting that the impact of high-quality captions is not dependent on the model variation. Additionally, compared to models like ShareGPT-4V, our model demonstrates superior performance across most VQA benchmarks. This improvement indicates that the captions generated by our EDC method provide richer and more comprehensive information. The model trained on EDC-1M demonstrates exceptional performance on datasets such as VQAv2 [16] and Ai2D [21], highlighting that integrating detection models into EDC significantly enriches the diversity of objects in the generated captions, thereby boosting the model’s performance on object recognition benchmarks. Furthermore, incorporating relational attributes into EDC enriches the captions with detailed inter-object relationships, enhancing the model’s ability to capture and understand complex relationships, which further improves its performance on visual reasoning benchmarks such as GQA [18]. However, performance on tasks like TextVQA [55] is hindered by limitations in the open-source OCR model and the threshold settings; a high threshold restricts the model’s ability to capture finer textual details.
@@
   Annotation    OKVQA    GQA    ScienceQA    TextVQA    MMBench    $M$-Vet    SEED-Bench
@@
   + EDC    56.7    65.2    72.0    62.0    68.5    37.8    65.0
Figure 4: Visualization of EDC’s Attribute Fusion: Integrating Object and Relational Attributes for Richer, More Detailed Captions.
Large Multi-Modal Benchmarks. We further conduct the evaluation on five challenging large multi-modal benchmarks. The experimental results are shown in Table 4. It can be seen that both LLaVA-v1.5 and LLaVA-NeXT trained with EDC-1M achieve competitive performance on more complex LMM benchmarks, demonstrating that the improvements brought by EDC-1M are comprehensive. Our model outperforms both LLaVA-v1.5 [34] and LLaVA-NeXT [35] across all LMM benchmarks, demonstrating that high-quality image captions during pretraining significantly enhance model performance, even without altering the supervised fine-tuning (SFT) data. Compared to other image captioning methods, such as ShareGPT-4V [7], EDC-generated captions provide richer and more comprehensive scene information, significantly boosting model performance across most LMM benchmarks. However, due to the lack of Chinese data in EDC-1M, the model performs poorly on MMBench-CN [37]. This highlights the need for multilingual image captioning, which will be an area for future improvement in EDC. Additionally, the detection model in EDC introduces some noise, which may interfere with the model’s ability to accurately capture objects, leading to decreased performance on tasks like POPE [30]. Therefore, reducing this noise will be a key focus for future improvements in EDC.
@@ 4.4 Ablation Study
Comparing different annotation methods. We compared different image annotation methods, including human annotations, GenerateList LMM annotations, and our EDC. Specifically, we annotated 118K COCO images and conducted comparisons on LLaVA-v1.5 and LLaVA-NeXT. The experimental results are shown in Table 6. We found that the image captions generated by EDC improve LMM performance on downstream tasks more effectively than other annotation methods. Notably, compared to captions annotated by internVL2, EDC’s inclusion of object attributes significantly improves model performance on OKVQA and TextVQA tasks. Relational attributes enhance the model’s understanding of multi-object relationships, leading to a notable increase in GQA performance. Furthermore, superior results on complex LMM benchmarks like MM-Vet, MMBench, and SEED highlight that EDC-generated captions provide rich and comprehensive scene information.
Case Study. Figure 4 presents example captions generated by the EDC engine and the general MLLM LLaVA-NeXT 34B. It is evident that visual specialists within EDC capture detailed object and relational attributes, resulting in richer and more descriptive captions. For instance, in Figure 4(a), the OCR information highlights the precision of OCR specialist model, while the fine-grained model’s identification of specific aircraft types further enhances caption informativeness. Additionally, the relational attributes significantly enrich description by providing detailed spatial relationships between objects, underscoring the advantages of EDC in capturing comprehensive scene details. More visualization results are provided in the Appendix E.
@@ 5 Conclusion
In this paper, we introduce EDC, a novel image captioning engine that leverages off-the-shelf visual specialists to enhance caption quality and detail. Unlike existing methods that rely solely on LMMs or human annotations, EDC integrates visual specialists to simulate human perception and LLMs to emulate cognitive reasoning. This dual approach enables the generation of captions that are both visually rich and contextually precise. Extensive experiments show that incorporating visual specialists significantly improves model performance across various visual understanding and reasoning tasks, especially those requiring precise attribute and relationship recognition. Our work underscores the potential of specialized visual features in strengthening multimodal representations and offers a flexible framework for integrating additional visual expertise. We will release the EDC source code to support further research, allowing the community to seamlessly integrate additional visual specialists and enhance EDC’s captioning capabilities.
Appendix A More Details of EDC pipeline
A.1 The Specific Visual Expert Model
In EDC, we employ different visual expert models to obtain object-level and relation-level attribute labels. The main paper provides an overview of the process for obtaining different attributes and briefly introduces the types of Visual Specialists used. In Table 7, we will further elaborate on the off-the-shelf visual specialists employed in the EDC pipeline. Based on our experience, existing open-source OCR models exhibit suboptimal performance in accurately recognizing text in images. To enhance OCR accuracy, we leverage Baidu’s API  [^1] https://ai.baidu.com/tech/ocr/general, which offers superior text recognition capabilities. Additionally, for attributes such as logo, celebrity, and landmark recognition, the limited availability of labeled data poses a challenge. To mitigate this, we also utilize Baidu’s API  [^2] https://ai.baidu.com/tech/imagerecognition to better capture these attributes, ensuring more precise and comprehensive annotations. Meanwhile, as a simple and versatile framework, EDC allows users to flexibly select the most suitable visual specialist models based on their available resources.
A.2 The Detail of Relation Attributes
Relation Attributes play a crucial role in image captioning. Among them, HOI captures human-object interactions, while 2D Absolute Location describes an object’s position within the image. Since such information is relatively sparse, we retain all relations in the captions. In contrast, 2D/3D Relative Location encodes spatial relationships between objects, which can become abundant and redundant as the number of objects increases. To address this, we retain only one dominant relationship per image for 2D/3D Relative Location, prioritizing the one with the greatest relative distance to ensure accuracy.
A.3 The Prompt Templates of EDC
To effectively transform extracted visual attributes into natural language, EDC designs two structured prompt templates that guide the LLM to convert discrete visual signals into coherent and readable descriptions. The first template focuses on object-level attributes, integrating multiple properties of each individual object. The second focuses on relation-level attributes, summarizing the relationships among different objects. The entire process is designed to mimic human perception: first describing what each object is, and then explaining how these objects relate to each other within the scene.
For the object-level prompt template (as shown in Figure 5), the LLM uses structured attributes extracted by visual specialists—such as category, color, texture, emotion, fine-grained classification (e.g., replacing “dog” with “golden retriever”), OCR content, and depth—to check whether the reference caption already includes these details. If a specific attribute is missing, it will be added; if it conflicts with the extracted value, the information from the visual specialist takes priority. This process allows the model to merge all attributes into a richer region-level description without redundancy or hallucinated content. For example, the prompt explicitly defines conditional rules such as “If a fine-grained animal category is available, replace the coarse label with the specific species; otherwise, keep the original name.” Likewise, it guides the LLM to naturally embed OCR text, detected emotions, and visual context into the caption, producing region-level descriptions that are both accurate and fluent.
For the relation-level prompt template (as shown in Figure 6), the LLM incorporates the relationships among different objects into the overall image caption. Among these, the P2O relations are directly obtained from the HOI model, while other types of relations are derived from detection and depth models. Specifically, the detection model provides the count of objects and, based on their bounding boxes, determines both 2D absolute locations and 2D relative locations to describe spatial relationships on the image plane. The depth model, on the other hand, extracts a depth value for each object from its bounding box and the depth map. By comparing the depth values between objects, the model can infer their 3D spatial relationships, such as whether object A is in front of or behind object B. Moreover, incorporating the object-level captions further enriches the image caption with detailed visual information. This structured prompting process ensures that the model maintains spatial grounding while merging scattered regional details into one coherent and comprehensive image-level description.
In practice, both templates are implemented as single-turn structured prompts rather than multi-turn reasoning. Each prompt directly guides the LLM to generate captions based on the provided visual attributes and contextual cues, without requiring iterative dialogue. The object-level template focuses on enriching region descriptions with fine-grained visual details, while the relation-level template integrates spatial and interaction information to form a complete image-level caption. Together, these two prompts enable EDC to seamlessly connect perception-level attributes with natural language expression, producing captions that are more detailed, precise, and contextually grounded than those from conventional LMMs.
Table 7: The Specific Visual Expert Model of EDC.
  Detection Model   Depth Model   OCR Model   HoI Model   Emotion Model
  In-domain   Open world
  Group Detr [8]   LaMI-DETR [13]   Depth Anything V2 [66]   API1   RLIPv2 [70]    [49]
  Fine-Grained Model
  Animal   Plant   Aircrafts   Logo   Landmark   Food   Celebrity
  BioClip [57]   BioClip [57]   MMALNet [73]   API2   API2   PreNet [42]   API2
Appendix B Analysis of EDC datasets
The EDC dataset consists of two parts: EDC-1M, comprising 1 million diverse image-text pairs sampled from the Laion dataset [51], and EDC-118K, comprising 118,000 real image-text pairs from the COCO dataset [32]. Next, we will analyze the captions in EDC-118K and compare them with captions annotated by humans [9] and generic LMM models [10, 27].
B.1 The Caption Length
In general, the longer caption could convey more detailed visual content. We compared the caption length of EDC-118k with human annotations as well as captions generated by advanced MLLM models, InternVL2-26B [10] and LLaVA-Next-34B [27]. The results are summarized in Table 8. It was observed that human-generated captions were the shortest, as they typically focus on only the most salient objects. The InternVL2-26B could generate more longer captions, with approximately 106 tokens. The captions generated by LLaVA-Next-34B were the longest, averaging around 228 tokens, while EDC-118k produced captions with an average of 218 tokens, approximately 10 tokens fewer than those of LLaVA-Next-34B.
B.2 The Lexical Composition
We conducted a detailed analysis of the lexical composition of the captions to evaluate how effectively each model described the visual content. This analysis examined the variety, frequency, and distribution of different word categories, including nouns, verbs, adjectives, adverbs, numerals, and more. As shown in Fig. 7, EDC-118k contained the highest average number of lexical elements per sentence, demonstrating a more diverse and complex linguistic structure compared to other captions generated by other datasets. This richer composition indicates that EDC-118k was better equipped to deliver nuanced and detailed descriptions of visual scenes, using a wider range of grammatical constructs to convey more comprehensive information.
Figure 7: Comparison of lexical composition of the captions generated by different methods. The y-axis represents the average frequency of each class of lexical per sentence in the datasets.
Table 8: Comparison of the different caption datasets. The ”ATL” abbreviates the ”Average Token Length”. The token length is counted by the tokenlizer of Vicuna-v1.5
Cpationed by Image Source Samples ATL of Caption
Human COCO 118k 14.67
InternVL2-26B 105.80
LLaVA-NeXT-34B 227.68
EDC 217.71
B.3 The Word Clouds
In Fig. 8, we present word clouds for the captions generated by InternVL2-26B, LLaVA-Next-34B, and our EDC pipeline. These visualizations highlight the most frequently used words across the different captioning methods, providing an intuitive comparison of the lexical patterns and focus areas of each caption. By examining the word clouds, we observed that the captions generated by our EDC exhibited a notably diverse vocabulary. In particular, there was a significantly higher frequency of words describing the relative spatial relationships of objects, both in 2D and 3D space like ‘left side’, ‘right region’, ‘front’, and ‘behind’, compared to captions produced by other methods. This indicates that EDC not only captured a wider range of visual details but also excelled in conveying the spatial context of objects within the scenes, offering more comprehensive descriptions of the visual content.
Figure 8: Word Cloud of captions generated by InternVL2, LLaVA-Next and EDC.
Appendix C Training Details
We elaborate the training details and hyper-parameters used in our experiments for evaluating the effectiveness of EDC-1M generated by our EDC pipeline. The whole training step consists of three stages, as shown in Table. 9. During the pre-alignment stage, we exclusively train the projector, resulting in a more stable and consistent vision-language connection. In the pre-training phase, similar with ShareGPT4V [7], we unfreeze the Vision Encoder (VE) for the last 12 layers, the Language Model (LM), and the projector. Regarding the instruction tuning stage, we use the open-source LLaVA-mix-665K [34] and LLaVA-NeXT-data to fine-tune both the projector and language model of the LLaVA-v1.5 [34] and LLaVA-NeXT [35] models, respectively.
Table 9: Training details and hyper-parameters used in our experiments. ‘VE’ means the vision encoder of CLIP for the last 12 layers, and ‘LM’ refers to the language model.
   Hyper-parameter    Pre-aligning    Pre-training    Instruction Tuning
   Batch Size    256    256    128
   Learning Rate    2e-5    2e-5    2e-5
   LR Schedule    cosine decay
   LR Warmup Ratio    0.01    0.01    0.01
   Weight Decay    0    0    0
   Trainable Module    Projector    LLaVA-v1.5: Projector, VE, LM    LLaVA-v1.5: Projector, LM
   LLaVA-NeXT: Full Model    LLaVA-NeXT: Full Model
   Epoch    1    1    1
   Optimizer    AdamW
   DeepSpeed stage    3    3    3
   Dataset    EDC-1M    EDC-1M     LLaVA-v1.5: LLaVA-mix-665K
   LLaVA-NeXT: LLaVA-NeXT-data
Table 10: CircularEval multi-choice accuracy results on MMBench [37] dev set. We adopt the following abbreviations: LR for Logical Reasoning; AR for Attribute Reasoning; $R$ for Relation Reasoning; FP-C for Fine-grained Perception (Cross Instance); FP-S for Finegrained Perception (Single Instance); CP for Coarse Perception.
Annotation Method MMBench-CN MMBench
Overall LR AR $R$ FP-S FP-C CP Overall LR AR $R$ FP-S FP-C CP
InternVL2-26B [10] 56.9 28.8 58.8 59.1 54.9 44.1 74.0 64.8 35.6 68.3 57.4 70.3 52.4 77.4
LLaVA-NeXT-34B [27] 56.5 28.8 60.3 56.5 54.3 42.0 74.3 64.9 31.4 68.3 57.4 69.3 54.5 79.7
EDC 58.2 29.7 62.3 57.4 56.7 45.5 74.7 65.8 37.3 71.4 57.4 70.0 53.8 78.4
lightgray black InternVL2-26B [10] 58.8 31.4 60.3 51.3 56.0 51.0 78.4 66.7 36.4 71.4 59.1 66.9 62.9 80.1
LLaVA-NeXT-34B [27] 59.8 31.3 60.8 54.8 56.7 49.7 80.7 67.2 38.1 69.8 57.4 69.3 60.1 82.1
EDC 60.1 31.4 62.3 55.7 57.0 51.0 79.1 68.5 39.0 72.4 67.8 67.9 59.4 82.8
Appendix D Detalied Evaluation Results
To highlight the significant improvements EDC-generated image descriptions bring to model performance, we present a detailed evaluation of the MMbench results in Table C. These results demonstrate how high-quality image descriptions enhance the model’s capabilities, particularly in logical reasoning, attribute reasoning, and relational reasoning. This improvement is driven by EDC’s ability to accurately capture object relationships and detailed attributes within images, enabling more effective reasoning and a deeper understanding of object interactions and characteristics.
Moreover, EDC-generated descriptions achieve competitive performance in both fine-grained and coarse perception tasks, showcasing the effectiveness of integrating various visual experts to emulate manual annotation. This approach enriches visual information, resulting in higher accuracy and robustness in tasks like visual question answering and image comprehension.
Appendix E Visualizations on EDC
To visually demonstrate the quality of captions annotated by EDC, we compared them with captions generated by generic LMMs, such as InternVL2-26B and LLaVA-NeXT-34B. The visualization highlights the differences in caption quality, providing a clear comparison of EDC’s detailed and accurate descriptions against those produced by the LMMs. As shown in Figure 9, EDC captions consistently capture more nuanced object attributes, relationships, and contextual details, showcasing its superior annotation capabilities. This comparison underscores EDC’s effectiveness in generating high-quality captions that enhance downstream visual-language tasks.
Figure 9: A comparison of image captions generated by InternVL2-26B, LLaVA-Next-34B, and EDC. We highlight different types of information, including Object Attributes, OCR, HOI, 2D spatial relations and 3D spatial relations.